I need help on Addition and Subtraction of Linear Equations...The problem is
3x+y=17
-1x-x=23

Answer by stanbon(75887)

3x+y=17
-1x-x=23
-------------
I think you posted it wrong.
Look at your 2nd equation.
Cheers,
Stan H.